# SynBrain: Enhancing Visual-to-fMRI Synthesis via Probabilistic Representation Learning

### v2

@@ Abstract
Deciphering how visual stimuli are transformed into cortical responses is a fundamental challenge in computational neuroscience. This visual-to-neural mapping is inherently a one-to-many relationship, as identical visual inputs reliably evoke variable hemodynamic responses across trials, contexts, and subjects. However, existing deterministic methods struggle to simultaneously model this biological variability while capturing the underlying functional consistency that encodes stimulus information. To address these limitations, we propose SynBrain, a generative framework that simulates the transformation from visual semantics to neural responses in a probabilistic and biologically interpretable manner. SynBrain introduces two key components: (i) BrainVAE models neural representations as continuous probability distributions via probabilistic learning while maintaining functional consistency through visual semantic constraints; (ii) A Semantic-to-Neural Mapper acts as a semantic transmission pathway, projecting visual semantics into the neural response manifold to facilitate high-fidelity fMRI synthesis. Experimental results demonstrate that SynBrain surpasses state-of-the-art methods in subject-specific visual-to-fMRI encoding performance. Furthermore, SynBrain adapts efficiently to new subjects with few-shot data and synthesizes high-quality fMRI signals that are effective in improving data-limited fMRI-to-image decoding performance. Beyond that, SynBrain reveals functional consistency across trials and subjects, with synthesized signals capturing interpretable patterns shaped by biological neural variability. The code will be made publicly available.
@@ 1 Introduction
Figure 1: Overview of SynBrain for adaptive visual-to-fMRI synthesis and downstream decoding applications. SynBrain is trained on full fMRI recordings from a source subject and adapted to novel subjects using limited data. It generates semantically consistent neural responses that support brain functional analysis and enhance downstream decoding through synthetic data augmentation.
To address these limitations, we propose SynBrain, a generative framework that models fMRI responses as a semantic-conditioned, continuous probability distribution. SynBrain is designed to simulate the transformation from visual semantics to brain activity in a manner that is probabilistic and biologically interpretable. At the core of SynBrain is BrainVAE, a well-designed variational model that incorporates convolutional feature extractors and attention modules to mitigate unstable training dynamics observed in MLP-based variational architectures. Unlike deterministic autoencoders, BrainVAE learns a probabilistic latent space of fMRI responses explicitly conditioned on high-level visual semantics, allowing it to generate diverse fMRI signals that reflect neural variability while preserving functional consistency. Moreover, we incorporate Semantic-to-Neural (S2N) Mapper, a lightweight Transformer module that establishes a point-to-distribution mapping from fixed CLIP semantic embeddings to the probabilistic latent space of BrainVAE. S2N Mapper achieves one-step semantic-to-neural mapping for stable and semantically coherent fMRI synthesis, avoiding the distribution mismatch problems observed in previous multi-step diffusion-based method bao2025mindsimulator .
@@ 2.1 BrainVAE
Motivation and Architecture Comparison. MindSimulator bao2025mindsimulator has employed MLP-based autoencoders (MLP-AE) to map high-dimensional fMRI signals into a deterministic latent space. To explore benefits of variational modeling, we first incorporated a variational formulation into this architecture, forming a baseline termed MLP-VAE. However, we found that MLP-VAE suffers from unstable training, with collapsed reconstructions and divergence in MSE loss (Figure 3). This might be attributed to MLP’s lack of spatial inductive bias and token-wise independence. To tackle this issue, we propose BrainVAE, a well-designed variational architecture for stable and effective fMRI representation learning. As shown in Figure 3-Left, BrainVAE integrates convolutional layers to extract local voxel features and attention layers to capture long-range inter-voxel dependencies, leading to a smoother and more expressive latent space for enhanced fMRI synthesis. Figure 3-Right demonstrates that BrainVAE empirically possesses two clear advantages in fMRI signal synthesis: 1) Improved voxel-level reconstruction fidelity and faster convergence with lower validation MSE within fewer epochs; 2) Stronger semantic expressiveness captured in underlying neural patterns indicated by lower validation CLIP loss and higher cross-modal retrieval accuracies.
@@
$\mathcal{L}_{\mathrm{BrainVAE}}=\mathcal{L}_{\mathrm{MSE}}+\lambda_{\mathrm{KL}}\mathcal{L}_{\mathrm{KL}}+\lambda_{\mathrm{CLIP}}\mathcal{L}_{\mathrm{CLIP}}$ (4)
where we set $\lambda_{\mathrm{KL}}=0.001$ to softly regularize the latent distribution while preserving reconstruction quality, and $\lambda_{\mathrm{CLIP}}=1000$ to ensure fast convergence of semantic alignment within a few epochs.
@@ 2.2 S2N Mapper
Compared to diffusion-based alignment, like Diffusion Transformer dit used in MindEye mindeye and MindSimulator bao2025mindsimulator , our one-step mapping strategy eliminates the need for iterative denoising and avoids the training–inference distribution gap (see Appendix E) by operating entirely within a semantically structured latent space. Instead of starting from random noise, the S2N Mapper receives well-formed CLIP embeddings as input and directly predicts fMRI latent representations that lie within the target manifold. This one-step, point-to-distribution mapping eliminates the need for handcrafted priors, repeated sampling, or post-hoc averaging used in MindSimulator bao2025mindsimulator , facilitating a stable and computationally efficient visual-to-fMRI synthesis process (see Appendix E for details).
@@ 3 Experimental Settings
We conduct experiments on the Natural Scenes Dataset (NSD) nsd , a large-scale fMRI dataset in which 8 subjects viewed natural images from the COCO dataset COCO across approximately 40 hours of scanning. Following MindSimulator bao2025mindsimulator , we focus on 4 subjects (Sub-1, Sub-2, Sub-5, Sub-7) who completed all experimental sessions. For each subject, we use 9,000 unique images for training and evaluate on a shared set of 1,000 test images, each presented across 3 trials to account for response variability. See Appendix B for additional dataset details.
Implementation Details. We adopt the pretrained OpenCLIP ViT-bigG/14 model clip as a frozen visual encoder to extract semantic embeddings from visual stimuli. Our model is implemented and trained on 4 NVIDIA A100 GPUs (40GB memory per GPU), with training completed within 2 hours.
Training Settings. We investigate visual-to-fMRI synthesis under two distinct settings. Additionally, we explore leveraging the generated fMRI signals for the decoding task in a few-shot setting.
1) Subject-specific full-data (40-hour) training: In this setting, we train BrainVAE using the AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.999)$, a learning rate of $1\times 10^{-4}$, weight decay of $0.05$, and a batch size of $24$. We apply early stopping to prevent overfitting. The S2N Mapper is optimized with identical hyperparameters and trained for 50,000 steps.
2) Novel-subject few-shot (1-hour) adaptation: In this setting, we retain the same training protocol as for source subjects but adopt a parameter-efficient strategy. Specifically, we finetune the entire BrainVAE while updating only the MLP submodules in S2N Mapper’s Transformer architecture.
3) Data augmentation for few-shot fMRI-to-Image decoding: In this setting, we follow the few-shot experimental protocol of MindEye2, with one key modification: we augment the limited real fMRI data using synthetic fMRI signals generated from unseen images, thereby expanding the training set. This setting helps to validate the quality of synthesized fMRI signals.
@@
1) Voxel-Level Metrics: To quantify reconstruction fidelity, we compute three voxel-wise similarity metrics between synthesized fMRI signals and all original trials, and report the average scores across trials: mean squared error (MSE), Pearson correlation (Pearson), and cosine similarity (Cosine).
2) Semantic-Level Metrics: To assess semantic quality, we use MindEye2 mindeye2 , a pretrained diffusion-based fMRI-to-image decoder. We evaluate the semantic alignment between decoded and original images using four established metrics: Inception Score (Incep) inception3 , CLIP similarity (CLIP) clip , EfficientNet distance (Eff) eff , and SwAV distance (SwAV) swav .
3) Image Retrieval Accuracy: We measure how well fMRI signals support cross-modal retrieval by computing cosine similarity between fMRI embeddings and CLIP visual embeddings. We compare two signal sources: (i) raw fMRI (Raw) and (ii) synthetic fMRI signals (Syn) from SynBrain. Retrieval results are averaged over 30 sampled subsets of 300 candidate images.
Please refer to Appendix C for detailed descriptions of these metrics and all fMRI-to-image decoding evaluation metrics.
@@ 4.1 Main Results
We first evaluate the quality of fMRI synthesis under the subject-specific setting, where both training and evaluation are performed on the same individual. As shown in the top section of Table 1, SynBrain (Trials=1) significantly outperforms MindSimulator across all voxel-level and semantic-level metrics, despite using only one-shot generation. These results highlight the effectiveness of SynBrain in producing faithful and semantically consistent neural signals without the need for repeated sampling or handcrafted priors used in MindSimulator.
@@
Figure 4 illustrates that even repeated trials of the same visual stimulus elicit fMRI responses with noticeable variability in fine details. SynBrain effectively learns to abstract away such fine-grained fluctuations while capturing consistent semantic patterns across trials (see Appendix F for diverse yet semantically consistent synthesis for identical visual stimulus). This is further evidenced by the fMRI-to-image visualizations via pretrained MindEye2, which reflect high-level concepts aligned with the original visual stimuli (e.g., teddy bear, man, flower, bus) despite fine-grained differences between the synthesized and raw fMRI signals.
@@
Table 2 reports quantitative results across multiple fMRI-to-image reconstruction and retrieval metrics. Despite using the same 1-hour real dataset, models augmented with synthetic signals consistently outperform the baselines, except for the image retrieval performance. Notably, MindEye2+DA(1h) achieves a higher CLIP similarity (84.7% vs. 80.8%), better Inception Score (85.1 vs. 83.6), and stronger brain retrieval performance (82.0% vs. 77.6%). Interestingly, the benefit plateaus or slightly declines with more synthetic data, suggesting that moderate augmentation yields the best balance between diversity and quality. These findings demonstrate that our high-quality synthetic fMRI signals can serve as an effective form of data augmentation, improving fMRI-to-image decoding in few-shot setting without requiring time-consuming and laborious data collection procedures.
@@ Acknowledgments
Appendix A Related Work
Visual-to-fMRI Encoding.
Few-Shot fMRI-to-Visual Decoding.
The task of fMRI-to-visual decoding aims to reconstruct visual stimuli from recorded fMRI responses, with related neural decoding efforts also explored in the Electroencephalography (EEG) guo2025neuro ; lu2025unimind ; zhou2025csbrain and Magnetoencephalography (MEG) benchetrit2023brain ; yang2024mad domains. Recent approaches typically map fMRI signals into pretrained vision-language embedding spaces (e.g., CLIP) and decode them into images using powerful generative backbones mind-vis ; mai2023unibrain ; Takagi2023 ; mindeye . While subject-specific decoding has achieved impressive results, however, collecting high-quality fMRI data is expensive, time-consuming, and subject to strong individual variability—making large-scale subject-specific training impractical in many scenarios. To address this, recent studies have explored cross-subject few-shot adaptation, where models are adapted to novel individuals using only 1 hour of brain recordings mindeyev1 ; mindeyev2 ; mindbridge ; mindtuner .
Appendix B NSD Dataset
@@
Appendix C Evaluation Metric
C.1 fMRI-to-Image Decoding Model.
@@ Appendix C Evaluation Metric
C.2 Visual-to-fMRI Synthesis Metrics
@@
C.3 Few-shot fMRI-to-Image Decoding Metrics
@@
Same as Semantic-Level Metrics via fMRI-to-Image Decoding in Section C.2.
@@
Appendix D Network Architecture
D.1 BrainVAE Architecture
@@
The encoder design is detailed in Algorithm 2. The input fMRI signal $x\in\mathbb{R}^{1\times V_{s}}$ first passes through a 1D convolution layer with 128 output channels and kernel size 7, followed by adaptive max pooling to a fixed length of 8192. This allows the model to process inputs from different subjects without introducing subject-specific parameters.
@@ D.2 S2N Mapper Architecture
Figure 6: UMAP visualizations of distribution gap in DiT and fMRI-to-image visualizations. Left: Noisy fMRI representations (blue) used for DiT training still lie close to the original fMRI representations (orange), but far away from pure Gaussian noise (green) used for DiT inference, showing a clear distribution gap between training and testing stages in DiT. Right: SynBrain (Brain-S2N, one-step mapping started from original fMRI representations) produces more realistic and semantically consistent images compared to BrainVAE-DiT (multi-step denoising started from Gaussian noise).
@@
Appendix E Comparison between Diffusion Transformer and S2N Mapper
Distribution Mismatch in Diffusion Prior.

### v3

@@ Abstract
Deciphering how visual stimuli are transformed into cortical responses is a fundamental challenge in computational neuroscience. This visual-to-neural mapping is inherently a one-to-many relationship, as identical visual inputs reliably evoke variable hemodynamic responses across trials, contexts, and subjects. However, existing deterministic methods struggle to simultaneously model this biological variability while capturing the underlying functional consistency that encodes stimulus information. To address these limitations, we propose SynBrain, a generative framework that simulates the transformation from visual semantics to neural responses in a probabilistic and biologically interpretable manner. SynBrain introduces two key components: (i) BrainVAE models neural representations as continuous probability distributions via probabilistic learning while maintaining functional consistency through visual semantic constraints; (ii) A Semantic-to-Neural Mapper acts as a semantic transmission pathway, projecting visual semantics into the neural response manifold to facilitate high-fidelity fMRI synthesis. Experimental results demonstrate that SynBrain surpasses state-of-the-art methods in subject-specific visual-to-fMRI encoding performance. Furthermore, SynBrain adapts efficiently to new subjects with few-shot data and synthesizes high-quality fMRI signals that are effective in improving data-limited fMRI-to-image decoding performance. Beyond that, SynBrain reveals functional consistency across trials and subjects, with synthesized signals capturing interpretable patterns shaped by biological neural variability. Our code is available at https://github.com/MichaelMaiii/SynBrain.
Figure 1: Overview of SynBrain for subject-adaptive visual-to-fMRI synthesis and downstream decoding applications. SynBrain is trained on full fMRI recordings from a source subject and adapted to novel subjects using limited data. It generates semantically consistent neural responses that support brain functional analysis and enhance downstream decoding through synthetic data augmentation.
@@ 1 Introduction
To address these limitations, we propose SynBrain, a generative framework that models fMRI responses as a semantic-conditioned, continuous probability distribution. SynBrain is designed to simulate the transformation from visual semantics to brain activity in a manner that is probabilistic and biologically interpretable. At the core of SynBrain is BrainVAE, a well-designed variational model that incorporates convolutional feature extractors and attention modules to mitigate unstable training dynamics observed in MLP-based variational architectures. Unlike deterministic autoencoders, BrainVAE learns a probabilistic latent space of fMRI responses explicitly conditioned on high-level visual semantics, allowing it to generate diverse fMRI signals that reflect neural variability while preserving functional consistency. Moreover, we incorporate Semantic-to-Neural (S2N) Mapper, a lightweight Transformer module that establishes a point-to-distribution mapping from fixed CLIP semantic embeddings to the probabilistic latent space of BrainVAE. S2N Mapper achieves one-step semantic-to-neural mapping for stable and semantically coherent fMRI synthesis, avoiding the distribution mismatch problems observed in previous multi-step diffusion-based methods bao2025mindsimulator .
@@ 2.1 BrainVAE
Motivation and Architecture Comparison. MindSimulator bao2025mindsimulator has employed MLP-based autoencoders (MLP-AE) to map high-dimensional fMRI signals into a deterministic latent space. To explore the benefits of variational modeling, we first incorporated a variational formulation into this architecture, forming a baseline termed MLP-VAE. However, we found that MLP-VAE suffers from unstable training, with divergence in MSE loss (Figure 3). This might be attributed to MLP’s lack of spatial inductive bias and token-wise independence. To tackle this issue, we propose BrainVAE, a well-designed variational architecture for stable and effective fMRI representation learning. As shown in Figure 3-Left, BrainVAE integrates convolutional layers to extract local voxel features and attention layers to capture long-range inter-voxel dependencies, leading to a smoother and more expressive latent space for enhanced fMRI synthesis. Figure 3-Right demonstrates that BrainVAE empirically possesses two clear advantages in fMRI signal synthesis: 1) Improved voxel-level reconstruction fidelity and faster convergence with lower validation MSE within fewer epochs; 2) Stronger semantic expressiveness captured in underlying neural patterns indicated by lower validation CLIP loss and higher cross-modal retrieval accuracies.
@@
$\mathcal{L}_{\mathrm{BrainVAE}}=\mathcal{L}_{\mathrm{MSE}}+\lambda_{\mathrm{KL}}\mathcal{L}_{\mathrm{KL}}+\lambda_{\mathrm{CLIP}}\mathcal{L}_{\mathrm{CLIP}}.$ (4)
Here we set $\lambda_{\mathrm{KL}}=0.001$ to softly regularize the latent distribution while preserving reconstruction quality, and $\lambda_{\mathrm{CLIP}}=1000$ to ensure fast convergence of semantic alignment within a few epochs.
@@ 2.2 S2N Mapper
Compared to diffusion-based alignment, like Diffusion Prior dalle2 used in MindEye mindeye and Diffusion Transformer dit used in MindSimulator bao2025mindsimulator , our one-step mapping strategy eliminates the need for iterative denoising and avoids the training–inference distribution gap (see Appendix D) by operating entirely within a semantically structured latent space. Instead of starting from random noise, the S2N Mapper receives well-formed CLIP embeddings as input and directly predicts fMRI latent representations that lie within the target manifold. This one-step, point-to-distribution mapping eliminates the need for handcrafted priors, repeated sampling, or post-hoc averaging used in MindSimulator bao2025mindsimulator , facilitating a stable and computationally efficient visual-to-fMRI synthesis process (see Appendix D for details).
@@ 3 Experimental Settings
We conduct experiments on the Natural Scenes Dataset (NSD) nsd , a large-scale fMRI dataset in which 8 subjects viewed natural images from the COCO dataset COCO across approximately 40 hours of scanning. Following MindSimulator bao2025mindsimulator , we focus on 4 subjects (Sub-1, Sub-2, Sub-5, Sub-7) who completed all experimental sessions. For each subject, we use 9,000 unique images for training and evaluate on a shared set of 1,000 test images, each presented across 3 trials to account for response variability. See Appendix A for additional dataset details.
Implementation Details. We adopt the pretrained OpenCLIP ViT-bigG/14 model clip as a frozen visual encoder to extract semantic embeddings from visual stimuli. Our model is implemented and trained on 4 NVIDIA A100 GPUs (40GB memory per GPU), with training completed within 2 hours. Note that a single NVIDIA A100 GPU is sufficient to train a variant of SynBrain (dd=1024) with trivial performance degradation (see Appendix Table 4 for details).
Training Settings. We explore visual-to-fMRI synthesis across two settings and assess whether the generated fMRI signals can effectively support decoding in few-shot scenarios.
i) Subject-specific full-data (40-hour) training: In this setting, we train BrainVAE using the AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.999)$, a learning rate of $1\times 10^{-4}$, weight decay of $0.05$, and a batch size of $24$. We apply early stopping to prevent overfitting. The S2N Mapper is optimized with identical hyperparameters and trained for 50,000 steps.
ii) Novel-subject few-shot (1-hour) adaptation: In this setting, we retain the same training protocol as for source subjects but adopt a parameter-efficient strategy. Specifically, we finetune the entire BrainVAE while updating only the MLP submodules in S2N Mapper’s Transformer architecture.
iii) Data augmentation for few-shot fMRI-to-Image decoding: In this setting, we follow the few-shot experimental protocol of MindEye2, with one key modification: we augment the limited real fMRI data using synthetic fMRI signals generated from unseen images, thereby expanding the training set. This setting helps to validate the quality of synthesized fMRI signals.
@@
i) Voxel-Level Metrics: To quantify reconstruction fidelity, we compute three voxel-wise similarity metrics between synthesized fMRI signals and all original trials, and report the average scores across trials: mean squared error (MSE), Pearson correlation (Pearson), and cosine similarity (Cosine).
ii) Semantic-Level Metrics: To assess semantic quality, we use MindEye2 mindeye2 , a pretrained diffusion-based fMRI-to-image decoder. We evaluate the semantic alignment between decoded and original images using four established metrics: Inception Score (Incep) inception3 , CLIP similarity (CLIP) clip , EfficientNet distance (Eff) eff , and SwAV distance (SwAV) swav .
iii) Image Retrieval Accuracy: We measure how well fMRI signals retain semantic information by computing cosine similarity between fMRI embeddings and CLIP visual embeddings. We compare two signal sources: (i) raw fMRI (Raw) and (ii) synthetic fMRI signals (Syn) from SynBrain. Retrieval results are averaged over 30 sampled subsets of 300 candidate images.
Please refer to Appendix B for detailed descriptions of these metrics and other fMRI-to-image decoding evaluation metrics.
@@ 4.1 Main Results
We first evaluate the quality of fMRI synthesis under the subject-specific setting, where both training and evaluation are performed on the same individual. As shown in the top section of Table 1, SynBrain (Trials=1) significantly outperforms MindSimulator bao2025mindsimulator across all voxel-level and semantic-level metrics, despite using only one-shot generation. These results highlight the effectiveness of SynBrain in producing faithful and semantically consistent neural signals without the need for repeated sampling or handcrafted priors used in MindSimulator.
@@
Figure 4 illustrates that even repeated trials of the same visual stimulus elicit fMRI responses with noticeable variability in fine details. SynBrain effectively learns to abstract away such fine-grained fluctuations while capturing consistent semantic patterns across trials (see Appendix E for diverse yet semantically consistent synthesis for identical visual stimulus). This is further evidenced by the fMRI-to-image visualizations via pretrained MindEye2, which reflect high-level concepts aligned with the original visual stimuli (e.g., teddy bear, man, flower, bus) despite fine-grained differences between the synthesized and raw fMRI signals.
@@
Table 2 reports quantitative results across multiple fMRI-to-image reconstruction and retrieval metrics. Despite using the same 1-hour real dataset, models augmented with synthetic signals consistently outperform the baselines, except for the image retrieval performance. Notably, MindEye2+DA(1h) achieves a higher CLIP similarity (84.7% vs. 80.8%), better Inception Score (85.1 vs. 83.6), and stronger brain retrieval performance (82.0% vs. 77.6%). Interestingly, the benefit plateaus or slightly declines with more synthetic data, suggesting that moderate augmentation yields the best balance between diversity and quality. These findings demonstrate that our high-quality synthetic fMRI signals can serve as an effective form of data augmentation, improving fMRI-to-image decoding in few-shot settings without requiring time-consuming and laborious data collection procedures.
@@ 4.3 Brain Functional Analysis
5 Related Work
Visual-to-fMRI Encoding.
Few-Shot fMRI-to-Visual Decoding.
The task of fMRI-to-visual decoding aims to reconstruct visual stimuli from recorded fMRI responses, with related neural decoding efforts also explored in the Electroencephalography (EEG) guo2025neuro ; lu2025unimind ; zhou2025csbrain ; wu2025adabrain and Magnetoencephalography (MEG) benchetrit2023brain ; yang2024mad domains. Recent approaches typically map fMRI signals into pretrained vision-language embedding spaces (e.g., CLIP) and decode them into images using powerful generative backbones mind-vis ; mai2023unibrain ; Takagi2023 ; mindeye . While subject-specific decoding has achieved impressive results, however, collecting high-quality fMRI data is expensive, time-consuming, and subject to strong individual variability—making large-scale subject-specific training impractical in many scenarios. To address this, recent studies have explored cross-subject few-shot adaptation, where models are adapted to novel individuals using only 1 hour of brain recordings mindeyev1 ; mindeyev2 ; mindbridge ; mindtuner .
@@ Acknowledgments
Appendix A NSD Dataset
@@
Appendix B Evaluation Metric
B.1 fMRI-to-Image Decoding Model.
@@ Appendix B Evaluation Metric
B.2 Visual-to-fMRI Synthesis Metrics
@@
B.3 Few-shot fMRI-to-Image Decoding Metrics
@@
Same as Semantic-Level Metrics via fMRI-to-Image Decoding in Section B.2.
@@
Appendix C Network Architecture
C.1 BrainVAE Architecture
@@
The encoder design is detailed in Algorithm 2. The input fMRI signal $x\in\mathbb{R}^{1\times V_{s}}$ first passes through a 1D convolution layer with 128 output channels and a kernel size of 1, followed by adaptive max pooling to a fixed length of 8192. This allows the model to process inputs from different subjects without introducing subject-specific parameters.
@@ C.2 S2N Mapper Architecture
Figure 6: UMAP visualizations of distribution gap in DiT and fMRI-to-image visualizations. Left: Noisy fMRI representations (blue) used for DiT training still lie close to the original fMRI representations (orange), but far away from pure Gaussian noise (green) used for DiT inference, showing a clear distribution gap between training and testing stages in DiT. Right: SynBrain (BrainVAE-S2N, one-step mapping started from original fMRI representations) produces more realistic and semantically consistent images compared to BrainVAE-DiT (multi-step denoising started from Gaussian noise).
@@
Appendix D Comparison between Diffusion Transformer and S2N Mapper
Distribution Mismatch in Diffusion Transformer.